OpenStudy (gabylovesyou):

On the graph of the equation 3x + 2y = 18, what is the value of the y-intercept? −9 −6 6 9

OpenStudy (anonymous):

9

OpenStudy (anonymous):

Convert to slope-intercept form. y = mx + b 3x + 2y = 18 First subtract 3x to both sides: 2y = -3x + 18 Now divide 2 to both sides.

OpenStudy (anonymous):

it is 9 do you want an explanation?

OpenStudy (anonymous):

TOLD U

OpenStudy (anonymous):

@ImAmWavy Please do not give answers, it's against OpenStudy's Code-of-Conduct. We are here to guide the user to the answer, so we need his participation. http://openstudy.com/code-of-conduct

OpenStudy (luigi0210):

Don't give away direct answers please :P And the y-intercept in short, is when x=0.

OpenStudy (gabylovesyou):

@iGreen. -3x + 18 / 2y ?

OpenStudy (anonymous):

No..we're dividing ONLY 2..not '2y'..

OpenStudy (gabylovesyou):

so y = -3x + 18 / 2 ?

OpenStudy (anonymous):

Oh that's right. @Luigi0210 ..we could also plug in 0 for 'x'..which sounds easier.

OpenStudy (anonymous):

y = -3x/2 + 18/2 is more like it..but we just write -3x/2 as -3/2x..so what's 18 / 2?

OpenStudy (gabylovesyou):

9

OpenStudy (anonymous):

Yes, so we have: y = -3/2x + 9 Now it's in the form of y = mx + b, where 'b' is the y-intercept, so what's the y-intercept here?

OpenStudy (gabylovesyou):

9

OpenStudy (gabylovesyou):

Thank you

OpenStudy (anonymous):

Yep, you got it.

OpenStudy (anonymous):

No

OpenStudy (anonymous):

*Np